OpenStudy (anonymous):

write a rule for the sequence -1, 2, 5, 8, 11, 14

OpenStudy (anonymous):

This is an arithmetic sequence, because each term is 3 greater than the term before it.

OpenStudy (anonymous):

rule for an arithmetic sequence is: a + (n-1)d a = first term d = difference In this sequence, the first term is -1, and the difference is 3. So the general term will be: -1 + (n-1)(3) = -1 + 3n - 3 = = -4 + 3n

OpenStudy (anonymous):

keep in mind, to find the common difference, take any terms, say the second and third term, and subtract the third term - second term... so 5 - 2 = 3.

OpenStudy (anonymous):

You're welcome.